# Exact model reduction for conditional quantum dynamics

### v1

@@ Abstract
Leveraging an algebraic approach built on minimal realizations and conditional expectations in quantum probability, we propose a method to reduce the dimension of quantum filters while maintaining the correct distributions on the outcomes of interest. The method is presented for general quantum systems whose dynamics depend on measurement outcomes, hinges on a system-theoretic observability analysis, and is tested on prototypical examples.
@@ I Introduction
Quantum stochastic dynamical models are key to the theory and practice of real-time feedback in quantum experiments [1], and at the same time represent one of its key limitations. In fact, despite their successful implementation in (relatively) small systems, the real-time integration of filtering equation poses significant limits to the effective bandwidth of the control. For this reasons, various methods have been considered to reduced the computational burden for these models, including projection filters and approximate reductions based on adiabatic-type limits [2, 3, 4, 5]. In this work, we take a system-theoretic route to the construction of stochastic quantum models in discrete time that are able to exactly reproduce the output and measurement statistics for a wide class of quantum systems, including discrete-time filtering and feedback equations [6, 7, 8], quantum trajectories [9], quantum walks including measurements [10], [11] and partially-observed quantum Markov process [12], while allowing for imperfect measurements and/or un-monitored environments.
@@
Our approach builds on the seminal work [13], (later generalized for quantum random walks by [14] using different tools), where stochastic analogues of reachable and non-observable spaces are used to characterize equivalent hidden-Markov processes. Exploiting the ideas developed in [15, 16] for unconditional dynamics, the method leverages algebraic quantum probability tools, in particular conditional expectations, to allow for reduced conditional models that generate the same measurement distribution, reproduce the average of observables of interests, and do so while maintaining the same structure and physical constraints of the original one. In general, existing reduction approaches cannot guarantee that the reduced evolution is completely positive (CP) and trace-preserving (TP) [17]. We thus show that the algebraic framework introduced in [16] is general and can be adapted to many cases of interest. Comparing the reduced models with and without conditioning, we highlight some key differences between classical and quantum processes: in some case quantum conditioning can further reduce the minimal realization, while in the classical case this is never possible. It is worth remarking that the same approach can be adapted to derive reduced models from known initial conditions of interest, i.e. exploiting reachable reductions. This problem is less natural for filtering dynamics, and requires the introduction and use of modified algebraic structures that go beyond the scope of the present work [16].
@@
In Section II we introduce the class of models of interest. Notice that some of the assumptions introduced in Section II for simplicity can be relaxed. For example, one could consider the case of multiple measurement apparatuses and evolution maps among which one can choose which one to measure or apply. Section III introduces the non-observable subspace and the optimal linear reduction. Similar ideas have been introduced, with different aims in [18]. Section IV illustrates how to extend a minimal observable representation of the dynamics to a CPTP one. Section V is devoted to examples.
@@ II Model and problem definition
In this paper, we consider finite-dimensional Hilbert spaces $\mathcal{H}\simeq\mathbb{C}^{n}$. The algebra of linear operators (closed with respect to standard matrix product) acting on $\mathcal{H}$ is denoted with $\mathfrak{B}(\mathcal{H})\simeq\mathbb{C}^{n\times n}$, and the set of denity operators as $\mathfrak{D}(\mathcal{H})=\{\rho|\rho\geq 0,\,{\rm tr}(\rho)=1\}$.
@@ II-A Dynamics of interest
The models we consider are aimed to cover a wide range of stochastic dynamics of practical interest such as complex quantum circuits with measurements [19, 20], or quantum experiments with repeated measurements [6, 21, 22]. More specifically, we want to be able to describe systems that evolve for some time, undergo a generalized measurement (possibly imperfect), and experience further dynamics conditioned on the outcome of the measurement.
A detailed account of indirect measurement and CPTP dynamics can be found in [23, 24]. Let us assume that we are able to perform a generalized measurement with outcomes $k\in\Omega$ described by the operators $\{M_{k}\}$ such that $\sum_{k}M_{k}^{\dagger}M_{k}=\mathbb{1}$. Some of the scenarios of interest for this work include:
$\bullet$ CPTP dynamics followed by generalized measurement: assume that the system evolves through a CPTP dynamical map $\mathcal{E}(\cdot)$ (e.g. evolution in contact with a Markov bath) and after each evolution the generalized measurement is performed. In this case, the effects of conditioning, given the measurement outcome $k$, are described by a set of CP maps $\{M_{k}\mathcal{E}(\cdot)M_{k}^{\dagger}\}$. $\bullet$ Imperfect measurement [21]: Assume now that the measurement apparatus correctly registers the measurement outcome $k$ with (known) probability $p_{k,k},$ while with probability $p_{k,j}$ the actual outcome was $j,$ with $\sum_{j}p_{k,j}=1$. Then the conditioning of the state given the registered outcome $k$ is given by a set of CP maps $\{\sum_{j}p_{k,j}M_{j}\cdot M_{j}^{\dagger}\}$.
@@
Let us consider a quantum system defined over a Hilbert space $\mathcal{H}$. Let us assume to be able to perform a measurement on the system at hand (either projective, generalized, or imperfect). Let $\Omega$ denote the set of outcomes of the measurement and let $\mathbb{P}_{\rho}[\mathcal{M}=k]$ be the probability of observing the outcome $k\in\Omega$ when performing the measurement on the system in the state $\rho$. We define a quantum instrument [25] as a set of CP super operators $\{\mathcal{M}_{k}\}_{k\in\Omega}$, one per every outcome in $\Omega$, such that $\sum_{k\in\Omega}\mathcal{M}_{k}^{\dagger}(\mathbb{1})=\mathbb{1}$.
Assume now that the system is prepared in a state $\rho$. The outcome $k\in\Omega$, is observed with probability
@@
Notice that the normalization factor at the denominator of the state update (2) makes the dynamics of the process nonlinear. Linear dynamics can be recovered by considering the un-normalized state $\tilde{\rho}|_{\mathcal{M}=k}=\mathcal{M}_{k}(\rho).$ We can further notice that, the probability of the outcome coincides with the trace of the un-normalized state, i.e. $\mathbb{P}_{\rho}[\mathcal{M}=k]={\rm tr}[\tilde{\rho}|_{\mathcal{M}=k}]$, and that the normalized state $\rho|_{\mathcal{M}=k}$ can still be retrieved by a posteriori re-normalization, i.e. $\rho|_{\mathcal{M}=k}={\tilde{\rho}|_{\mathcal{M}=k}}/{{\rm tr}[\mathcal{M}_{k}(\rho)]}.$
@@ II-B Quantum Conditional Evolutions
We now introduce a class of discrete-time quantum dynamical models that includes the effects of conditioning. In the following, we assume: 1) To have a quantum instrument, described by a set of CP maps $\{\mathcal{M}_{k}\}_{k\in\Omega}$ such that $\sum_{k\in\Omega}\mathcal{M}_{k}^{{\dagger}}(\mathbb{1})=\mathbb{1}$, where $\Omega$ is the set of possible outcomes; 2) That a measurement is performed at each time step $t=0,1,2,\dots$. We will thus indicate the outcome observed at time $t$ with the subscript $t$, i.e. $\mathcal{M}(t)=k_{t}$ indicates the measurement at time $t$ lead to the outcome $k_{t}\in\Omega$. Moreover, we denote an ordered sequence of conditional dynamics as $\mathcal{M}_{k_{0:t}}$ and of outcomes as $k_{0:t}:=k_{0},k_{1},\dots,k_{k}$ and thus
@@
3) We further assume to only be interested in reproducing the expectation value of a set of observables $\{O_{i}\}$ containing the identity, i.e. $\mathbb{1}\in\{O_{i}\}$. We compactly represent the set of observables of interest $\{O_{i}\}$ via a linear output map $\mathcal{C}:\mathfrak{B}(\mathcal{H})\to\mathscr{Y}$ where $\mathscr{Y}$ is a vector space of dimension $\dim(\mathscr{Y})=|\{O_{i}\}|$. This map is constructed by fixing a base $\{E_{i}\}$ for $\mathscr{Y}$ and imposing $\mathcal{C}(\cdot)=\sum_{i}E_{i}{\rm tr}[O_{i}\cdot]$.
@@
In exactly the same manner, it is possible to retrieve the evolution of the expectation value of an observable $O$ conditioned on the outcomes $k_{0:t}$, i.e. $O(t+1|\mathcal{M}(0:t)=k_{0:t})\right>$, starting from the un-normalized state via a posteriori re-normalization, ${\rm tr}[O\rho(t+1|\mathcal{M}(0:t)=k_{0:t}))]={{\rm tr}[O\tilde{\rho}(t+1,\mathcal{M}(0:t)=k_{0:t}))]}/{\mathbb{P}_{\rho_{0}}[\mathcal{M}({0:t})=k_{0:t}]}.$ Thus, in order to reproduce the evolution of the expectation value of an observable $O$ we simply need to keep track of the evolution of the two quantities ${\rm tr}[O\tilde{\rho}(t)]$ and ${\rm tr}[\tilde{\rho}(t)]$.
@@ III Observability and Linear Reduction
In this section, we will extend the work of [13] to the quantum framework. We shall start by defining the minimal subspace that supports the dynamics. This is done by adapting well-known results of control system theory [26, 27] to the case of switching systems.
@@
$\mathscr{N}$ =$=\left\{X|\mathcal{C}[X]=0\text{ and }\mathcal{C}\left[\mathcal{M}_{k_{0:t}}[X]\right]=0,\,\forall k_{0:t},t\geq 0\right\}$
@@
Notice that the orthogonal subspace to $\mathscr{N}$ takes the form
$\mathscr{N}^{\perp}={\rm span}\left\{O_{i},\mathcal{M}_{k_{0:t}}^{{\dagger}}(O_{i}),\,\forall k_{0:t},\,t\geq 0,\,\forall i\right\}.$ (5)
Moreover, one can also extend the results of [13, Lemma 3] to prove that the sequences $k_{0:t}$ can be limited to sequences of length less or equal to $\dim(\mathcal{H})^{2}$. Finally, we can notice that, because we assumed that $\mathbb{1}\in\{O_{i}\}$, we have that $\mathbb{1}\in\mathscr{N}^{\perp}$.
@@ III-A Linear model reduction
The proof of this follows from [15, Theorem 4]. To obtain the linear reduced model one can then find two factors of $\Pi_{\mathscr{N}^{\perp}}$, $\mathcal{R}_{L}:\mathfrak{B}(\mathcal{H})\to\mathbb{C}^{q}$ and $\mathcal{J}_{L}:\mathbb{C}^{q}\to\mathscr{N}^{\perp}$ with $q=\dim(\mathscr{N}^{\perp})$, such that $\Pi_{\mathscr{N}^{\perp}}=\mathcal{J}\mathcal{R}$ and $\mathcal{R}\mathcal{J}=\mathbb{1}_{q}$. With these, one can construct a set of $q\times q$ matrices $\{A_{k}=\mathcal{R}_{L}\mathcal{M}_{k}\mathcal{J}_{L}\}$ and an output matrix $C=\mathcal{C}\mathcal{J}_{L}$ such that the reduced model
@@ IV Quantum model reduction
IV-A ***-Algebras and conditional expectations
We here briefly review the main results from quantum probability theory, finite-dimensional operator algebras, and conditional expectations that are necessary for this work. more details on the matter can be found in [16, 23, 28, 29].
@@
$U^{\dagger}\mathscr{A}U=\bigoplus_{k}\mathfrak{B}(\mathcal{H}_{S,k})\otimes\mathbb{1}_{S,k}.$
An algebra with such a structure is thus isomorphic to a algebra $\widecheck{\mathscr{A}}=\bigoplus_{k}\mathfrak{B}(\mathcal{H}_{S,k}),$ which admits representation in a smaller Hilbert space. Furthermore, any unital algebra admits a CPTP and unital orthogonal projector $\mathbb{E}|_{\mathscr{A}}:\mathfrak{B}(\mathcal{H})\to\mathscr{A}$, i.e. $\mathbb{E}|_{\mathscr{A}}^{\dagger}=\mathbb{E}|_{\mathscr{A}}$, $\mathbb{E}|_{\mathscr{A}}(\mathbb{1})=\mathbb{1}$. Such a map is a conditional expectation (that preserves the completely mixed state $\mathbb{1}/n$) [23]. Conditional expectations can be factorized into two CP factors, which we denote with $\mathcal{R}:\mathfrak{B}(\mathcal{H})\to\widecheck{\mathscr{A}}$ and $\mathcal{J}:\widecheck{\mathscr{A}}\to\mathscr{A}$ such that $\mathcal{R}\mathcal{J}=\mathcal{I}_{\widecheck{\mathscr{A}}}$ the identity super operator over the algebra $\widecheck{\mathscr{A}}$, and $\mathcal{J}\mathcal{R}=\mathbb{E}|_{\mathscr{A}}$. Given the Wedderburn decomposition of $\mathscr{A}$, the conditional expectation and its CPTP factorization can be easily constructed [28, 16]. In the case of orthogonal conditional expectations $\mathbb{E}|_{\mathscr{A}}$ gets factorized as $\mathbb{E}|_{\mathscr{A}}=\mathcal{J}\circ\mathcal{R}$ where:

### v2

# Exact model reduction for discrete-time conditional quantum dynamics

@@ Abstract
Leveraging an algebraic approach built on minimal realizations and conditional expectations in quantum probability, we propose a method to reduce the dimension of quantum filters in discrete-time, while maintaining the correct distributions on the measurement outcomes and the expectations of some relevant observable. The method is presented for general quantum systems whose dynamics depend on measurement outcomes, hinges on a system-theoretic observability analysis, and is tested on prototypical examples.
@@ I Introduction
Quantum stochastic dynamical models are key to the theory and practice of real-time feedback in quantum experiments [1], and at the same time represent one of its key limitations. In fact, despite their successful implementation in (relatively) small systems, the real-time integration of filtering equations poses significant limits to the effective bandwidth of the control. For these reasons, various methods have been considered to reduce the computational burden of these models, including projection filters and approximate reductions based on adiabatic-type limits [2, 3]. In this work, we construct stochastic quantum models in discrete time that are able to exactly reproduce the output and measurement statistics for a wide class of quantum systems, including discrete-time filtering and feedback equations [4, 5, 6], quantum trajectories [7], quantum walks including measurements [8], [9] and partially observed quantum Markov process [10], while allowing for imperfect measurements and/or un-monitored environments.
@@
Our approach builds on the ideas of [11], (see [12] for a generalization to quantum walks), where stochastic analogues of reachable and non-observable spaces are used to characterize equivalent hidden-Markov processes. Exploiting the ideas developed in [13, 14] for unconditional dynamics, the method leverages algebraic quantum probability tools, in particular conditional expectations, to allow for reduced conditional models that generate the same measurement distribution, reproduce the average of observables of interest, and do so while maintaining the same structure and physical constraints as the original model. In general, existing reduction approaches cannot guarantee that the reduced evolution is completely positive (CP) and trace-preserving (TP) [15]. Therefore, we show that the algebraic framework introduced in [14] is general and can be adapted to many cases of interest. Comparing the reduced models with and without conditioning, we highlight some key differences between classical and quantum processes: in some cases quantum conditioning can further reduce the minimal realization, while in the classical case this is never possible. It is worth remarking that the same approach can be adapted to derive reduced models from known initial conditions of interest, i.e. exploiting reachable reductions. This problem is less natural for filtering dynamics and requires the introduction and use of modified algebraic structures that go beyond the scope of the present work [14].
@@
In Section II we introduce the class of models of interest. Note that some of the assumptions introduced in Section II for simplicity can be relaxed. For example, one could consider the case of multiple measurement apparatuses and evolution maps, among which one can choose which one to measure or apply. Section III introduces the non-observable subspace and the optimal linear reduction. Similar ideas have been introduced, with different objectives in [16]. Section IV illustrates how to extend a minimal observable representation of the dynamics to a CPTP one. Section V is devoted to examples.
@@ II Model and problem definition
In this paper, we consider finite-dimensional Hilbert spaces $\mathcal{H}\simeq\mathbb{C}^{n}$. The algebra of linear operators (closed with respect to the standard matrix product) acting on $\mathcal{H}$ is denoted by $\mathfrak{B}(\mathcal{H})\simeq\mathbb{C}^{n\times n}$, and the set of density operators by $\mathfrak{D}(\mathcal{H})=\{\rho|\rho\geq 0,\,{\rm tr}(\rho)=1\}$.
@@ II-A Dynamics of interest
Let us assume that we are able to perform a generalized measurement [17] with outcomes $k\in\Omega$ described by the operators $\{M_{k}\}$ such that $\sum_{k}M_{k}^{\dagger}M_{k}=\mathbb{1}$. Some of the scenarios of interest for this work include:
$\bullet$ CPTP dynamics followed by generalized measurement: assume that the system evolves through a CPTP dynamical map $\mathcal{E}(\cdot)$ (e.g. evolution in contact with a Markov bath) and after each evolution the generalized measurement is performed. In this case, the effects of conditioning, given the measurement outcome $k$, are described by a set of CP maps $\{M_{k}\mathcal{E}(\cdot)M_{k}^{\dagger}\}$.
$\bullet$ Imperfect measurement [6]: Assume now that the measurement apparatus correctly registers the measurement outcome $k$ with (known) probability $p_{k,k},$ while with probability $p_{k,j}$ the actual outcome was $j,$ with $\sum_{j}p_{k,j}=1$. Then the conditioning of the state given the registered outcome $k$ is given by a set of CP maps $\{\sum_{j}p_{k,j}M_{j}\cdot M_{j}^{\dagger}\}$.
@@
Let us consider a quantum system defined over a Hilbert space $\mathcal{H}$. Let us assume to be able to perform a measurement on the system at hand (either projective, generalized, or imperfect). Let $\Omega$ denote the set of measurement outcomes and let $\mathbb{P}_{\rho}[\mathcal{M}=k]$ be the probability of observing the outcome $k\in\Omega$ when performing the measurement on the system in the state $\rho$. A quantum instrument [17] is a set of CP super operators $\{\mathcal{M}_{k}\}_{k\in\Omega}$, one per every outcome in $\Omega$, such that $\sum_{k\in\Omega}\mathcal{M}_{k}^{\dagger}(\mathbb{1})=\mathbb{1}$. Assume that the system is prepared in a state $\rho$. The outcome $k\in\Omega$, is observed with probability
@@
Notice that the normalization factor at the denominator of the state update (2) makes the dynamics of the process non-linear. Linear dynamics can be recovered by considering the un-normalized state $\tilde{\rho}|_{\mathcal{M}=k}=\mathcal{M}_{k}(\rho).$ We can further notice that the probability of the outcome coincides with the trace of the un-normalized state, i.e. $\mathbb{P}_{\rho}[\mathcal{M}=k]={\rm tr}[\tilde{\rho}|_{\mathcal{M}=k}]$, and that the normalized state $\rho|_{\mathcal{M}=k}$ can still be retrieved by a posterior re-normalization, i.e. $\rho|_{\mathcal{M}=k}={\tilde{\rho}|_{\mathcal{M}=k}}/{{\rm tr}[\mathcal{M}_{k}(\rho)]}.$
@@ II-B Quantum Conditional Evolutions
We now introduce a class of discrete-time quantum dynamical models that includes the effects of conditioning. In the following, we assume: 1) To have a quantum instrument, described by a set of CP maps $\{\mathcal{M}_{k}\}_{k\in\Omega}$ such that $\sum_{k\in\Omega}\mathcal{M}_{k}^{{\dagger}}(\mathbb{1})=\mathbb{1}$, where $\Omega$ is the set of possible outcomes; 2) That a measurement is performed at each time step $t=0,1,2,\dots$. Therefore, we will indicate the outcome observed at time $t$ with the subscript $t$, i.e. $\mathcal{M}(t)=k_{t}$ indicates the measurement at time $t$ leads to the outcome $k_{t}\in\Omega$. Moreover, we denote an ordered sequence of conditional dynamics as $\mathcal{M}_{k_{0:t}}$ and of outcomes as $k_{0:t}:=k_{0},k_{1},\dots,k_{k}$ and thus
@@
3) We further assume to only be interested in reproducing the expectation value of a set of observables $\{O_{j}\}$ containing the identity, i.e. $\mathbb{1}\in\{O_{j}\}$. We compactly represent the set of observables of interest $\{O_{j}\}$ via a linear output map $\mathcal{C}:\mathfrak{B}(\mathcal{H})\to\mathscr{Y}$ where $\mathscr{Y}$ is a vector space of dimension $\dim(\mathscr{Y})=|\{O_{j}\}|$. This map is constructed by fixing a basis $\{E_{j}\}$ for $\mathscr{Y}$ and imposing $\mathcal{C}(\cdot)=\sum_{i}E_{i}{\rm tr}[O_{j}\cdot]$.
@@
In exactly the same manner, it is possible to retrieve the evolution of the expectation value of an observable $O$ conditioned on the outcomes $k_{0:t}$, i.e. $O(t+1|\mathcal{M}(0:t)=k_{0:t})\right>$, starting from the un-normalized state via a posterior re-normalization, ${\rm tr}[O\rho(t+1|\mathcal{M}(0:t)=k_{0:t}))]={{\rm tr}[O\tilde{\rho}(t+1,\mathcal{M}(0:t)=k_{0:t}))]}/{\mathbb{P}_{\rho_{0}}[\mathcal{M}({0:t})=k_{0:t}]}.$ Thus, in order to reproduce the evolution of the expectation value of an observable $O$ we simply need to keep track of the evolution of the two quantities ${\rm tr}[O\tilde{\rho}(t)]$ and ${\rm tr}[\tilde{\rho}(t)]$.
@@ III Observability and Linear Reduction
In this section, we will extend the work of [11] to the quantum framework. We shall start by defining the minimal subspace that supports the dynamics. This is done by adapting well-known results of control system theory [18] to the case of switching systems.
@@
$\mathscr{N}$ =$=\left\{X\in\mathfrak{B}(\mathcal{H})|\quad\mathcal{C}[X]=0\right.$
[…]
@@
Notice that the orthogonal subspace to $\mathscr{N}$ (w.r.t the standard Hilbert-Schmidt inner product ,$_{HS}$) takes the form
$\mathscr{N}^{\perp}={\rm span}\left\{O_{j},\mathcal{M}_{k_{0:t}}^{{\dagger}}(O_{j}),\,\forall k_{0:t},\,t\geq 0,\,\forall j\right\}.$ (5)
This can be derived from the definition by writing $\mathcal{C}(\mathcal{M}_{k_{0:t}}(X))=\sum_{i}E_{i}{\rm tr}[O_{j}\mathcal{M}_{k_{0:t}}(X)],$ interpreting the trace as the HS inner product and switching to the dual dynamics $\mathcal{M}_{k_{0:t}}^{\dagger}(O_{j}).$ It is then sufficient to notice that the latter operators are orthogonal to all the $X\in\mathscr{N}$. Moreover, one can also extend the results of [11, Lemma 3] to prove that the sequences $k_{0:t}$ can be limited to sequences of length less or equal to $\dim(\mathcal{H})^{2}$. Finally, we can notice that, because we assumed that $\mathbb{1}\in\{O_{j}\}$, we have that $\mathbb{1}\in\mathscr{N}^{\perp}$.
@@ III-A Linear model reduction
The proof of this follows from [13, Theorem 4] on the reducibility of switching dynamics. This can be applied to Proposition 2 by vectorizing the state and obtaining the corresponding matrix form of the superoperators. To obtain the linear reduced model one can then find two factors of $\Pi_{\mathscr{N}^{\perp}}$, $\mathcal{R}_{L}:\mathfrak{B}(\mathcal{H})\to\mathbb{C}^{q}$ and $\mathcal{J}_{L}:\mathbb{C}^{q}\to\mathscr{N}^{\perp}$ with $q=\dim(\mathscr{N}^{\perp})$, such that $\Pi_{\mathscr{N}^{\perp}}=\mathcal{J}\mathcal{R}$ and $\mathcal{R}\mathcal{J}=\mathbb{1}_{q}$. With these, one can construct a set of $q\times q$ matrices $\{A_{k}=\mathcal{R}_{L}\mathcal{M}_{k}\mathcal{J}_{L}\}$ and an output matrix $C=\mathcal{C}\mathcal{J}_{L}$ such that the reduced model
@@ IV Quantum model reduction
IV-A *-Algebras and conditional expectations
We here briefly review the main results from quantum probability theory, finite-dimensional operator algebras, and conditional expectations that are necessary for this work. More details on the matter can be found in [14, 19, 20].
@@
$U^{\dagger}\mathscr{A}U=\bigoplus_{k}\mathfrak{B}(\mathcal{H}_{S,k})\otimes\mathbb{1}_{F,k}.$
An algebra with such a structure is thus isomorphic to an algebra $\widecheck{\mathscr{A}}=\bigoplus_{k}\mathfrak{B}(\mathcal{H}_{S,k}),$ which admits representation in a smaller Hilbert space. Furthermore, any unital algebra admits a CPTP and unital orthogonal projector $\mathbb{E}|_{\mathscr{A}}:\mathfrak{B}(\mathcal{H})\to\mathscr{A}$, i.e. $\mathbb{E}|_{\mathscr{A}}^{\dagger}=\mathbb{E}|_{\mathscr{A}}$, $\mathbb{E}|_{\mathscr{A}}(\mathbb{1})=\mathbb{1}$ (where the adjoint here is taken w.r.t the inner product ,$_{HS}$). Such a map is a conditional expectation (that preserves the completely mixed state $\mathbb{1}/n$) [19]. Conditional expectations can be factorized into two CP factors, which we denote with $\mathcal{R}:\mathfrak{B}(\mathcal{H})\to\widecheck{\mathscr{A}}$ and $\mathcal{J}:\widecheck{\mathscr{A}}\to\mathscr{A}$ such that $\mathcal{R}\mathcal{J}=\mathcal{I}_{\widecheck{\mathscr{A}}}$ the identity super operator over the algebra $\widecheck{\mathscr{A}}$, and $\mathcal{J}\mathcal{R}=\mathbb{E}|_{\mathscr{A}}$. Given the Wedderburn decomposition of $\mathscr{A}$, the conditional expectation and its CPTP factorization can be easily constructed [20, 14]. In the case of orthogonal conditional expectations $\mathbb{E}|_{\mathscr{A}}$ gets factorized as $\mathbb{E}|_{\mathscr{A}}=\mathcal{J}\circ\mathcal{R}$ where: